OpenStudy (anonymous):

The hypotenuse of an isosceles right triangle is 2 cm. Find the measure of one of its legs

OpenStudy (anonymous):

Special case of Pythagorean theorem: \(a^2+b^2=c^2\) with \(a=b\). So, \(a^2+a^2=2a^2=c^2\). Your hypotenuse is \(c\); in other words, \(2 \text{ cm}\). Solve from here.